OpenStudy (anonymous):

The velocity of sound in air is given by the equation =20√(273+t) , where v is the velocity in meters per second and t is the temperature in degrees Celsius. Find the temperature when the velocity is 329 meters per second by graphing the equation. Round the answer to the nearest degree

OpenStudy (anonymous):

Your formula is \[v=20\sqrt{273+t}\] You have to find \(t\) for a given \(v\): \[329=20\sqrt{273+t}\] Do you know how to solve that?

OpenStudy (anonymous):

No how?

OpenStudy (anonymous):

Divide both sides by 20: \[\frac{329}{20}=\sqrt{273+t}\] Square both sides: \[\left(\frac{329}{20}\right)^2=273+t\] See where this is going?

OpenStudy (anonymous):

Umm yeah hold on let me try and solve ok?:)